What is a quartic function with only the two real zeros given?
x = -4 and x = -1

A. y = x^4 + 5x^3 + 5x^2 + 5x + 4
B. y = x^4 - 5x^3 - 5x^2 - 5x - 4
C. y = -x^4 + 5x^3 + 5x^2 + 5x + 4
D. y = x^4 + 5x^3 + 5x^2 + 5x - 5

Answer by ewatrrr(24785)

A. y = x^4 + 5x^3 + 5x^2 + 5x + 4
x = -1, y = 1 -5 + 5 - 5 + 4 = 0
x = -4, y = 0